OpenStudy (anonymous):

Determine Laplace transform of f(t)=10 sin(2t-30degrees) 1.F(s)=20/(s+2)^2+4 2.20cos30deg./s^2+4 3.-5s+17.32/s^2+4 4.8.66/s^2+4

OpenStudy (anonymous):

We need to work in radian mode, so \[L \left\{ 10 \sin(2t - \pi/6 \right\}\] is the function to work with. so \[ 2t - \pi/6 = 2(t - \pi/12), so we have L \left\{ \sin(2(t - \pi/12)) \right\}\],

OpenStudy (anonymous):

\[= 2/(s ^{2} + 4)\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

u didnt do any shifting at all